### v1

# ArchRAG: Attributed Community-based Hierarchical Retrieval-Augmented Generation

Abstract.
Retrieval-Augmented Generation (RAG) has proven effective in integrating external knowledge into large language models (LLMs) for question-answer (QA) tasks. The state-of-the-art RAG approaches often use the graph data as the external data since they capture the rich semantic information and link relationships between entities. However, existing graph-based RAG approaches cannot accurately identify the relevant information from the graph and also consume large numbers of tokens in the online retrieval process. To address these issues, we introduce a novel graph-based RAG approach, called Attributed Community-based Hierarchical RAG (ArchRAG), by augmenting the question using attributed communities, and also introducing a novel LLM-based hierarchical clustering method. To retrieve the most relevant information from the graph for the question, we build a novel hierarchical index structure for the attributed communities and develop an effective online retrieval method. Experimental results demonstrate that ArchRAG outperforms existing methods in terms of both accuracy and token cost.
Retrieval-Augmented Generation, Graph-based RAG
1. Introduction
Large Language Models (LLMs) like GPT-4  (Achiam et al., 2023), Qwen2.5  (Yang et al., 2024), and LLaMA3.1  (Dubey et al., 2024) have received tremendous attention from both industry and academia (Huang and Huang, 2024; Liu et al., 2024c; Wang et al., 2024b; Zheng et al., 2024; Li et al., 2023; Nie et al., 2024; Ghimire et al., 2024; Wang et al., 2024c). Despite their remarkable success in question-answering (QA) tasks, they may still generate wrong answers due to a lack of domain-specific, real-time updated, and proprietary knowledge outside their pre-training corpus (Peng et al., 2024). To enhance the trustworthiness and interpretability of LLMs, Retrieval-Augmented Generation (RAG) methods (Fan et al., 2024; Gao et al., 2023; Hu and Lu, 2024; Huang and Huang, 2024; Wu et al., 2024a; Yu et al., 2024; Zhao et al., 2024) have emerged as a core approach, which often retrieve relevant information from documents, relational data, and graph data to facilitate the QA tasks.
The state-of-the-art RAG approaches often use the graph data as the external data since they capture the rich semantic information and link relationships between entities. Given a question $Q$, the key idea of graph-based RAG is to retrieve relevant information (e.g., nodes, subgraphs, or textual information) from the graph, incorporate them with $Q$ as the prompt, and feed them into the LLM, as illustrated in Figure 1. As shown in the literature, a few recent graph-based RAG methods have emerged for both abstract question (Edge et al., 2024; Guo et al., 2024; Sarthi et al., 2024) and specific question (Guo et al., 2024; Sarthi et al., 2024; Wu et al., 2024b; Wang et al., 2024a; Li et al., 2024; Gutiérrez et al., 2024), where the former is more conceptual, encompassing broader topics, summaries, or overarching themes (e.g., “What are the potential impacts of LLMs on education?”), while the latter is detail-oriented and typically references specific entities within the graph (e.g., “Who has won the Turing Award in 2024?”) (Guo et al., 2024). Technically, the abstract questions need a comprehensive understanding to extract high-level insights, while the specific questions require multi-hop reasoning to connect discrete information. Table 1 summarizes the key characteristics of three representative graph-based RAG methods, i.e., GraphRAG (Edge et al., 2024), HippoRAG (Gutiérrez et al., 2024), and LightRAG (Guo et al., 2024).
Figure 1. The general workflow of graph-based RAG, which retrieves relevant information (e.g., nodes, subgraphs, or textual information) to facilitate the LLM generation.
Specifically, Microsoft first proposes a graph-based RAG system called GraphRAG (Edge et al., 2024), which first builds a knowledge graph (KG) by extracting entities and relationships from the external corpus, then employs the Leiden method (Traag et al., 2019) to detect communities and generates a summary for each community using LLMs. When answering abstract questions, it uses a Global Search method, which traverses each community and uses LLMs to retrieve relevant communities to answer the questions; for specific questions, it presents a Local Search method, which retrieves entities, identifies the relevant chunk texts and low-level communities to augments the response. LightRAG (Guo et al., 2024) addresses abstract questions by adding keywords to relations in the constructed KG, and during querying, it uses LLMs to extract keywords and searches for the most relevant entities and relationships, which are then combined with the corresponding chunks for generation. HippoRAG (Gutiérrez et al., 2024) employs LLMs to extract tuples from the corpus to build a KG and connect edges between similar entities in KG to improve the quality of the KG. To answer specific questions requiring multi-hop reasoning, HippoRAG utilizes the Personalized PageRank (PPR) (Haveliwala, 2002) to identify the most relevant entities from the KG and augment the question with chunks containing these selected entities for LLM generation.
Table 1. Comparison of representative RAG methods and our ArchRAG.
RAG method Retrieval method Question Attributed community Hierarchical index Retrieval filtering
[…]
While these approaches have achieved good performance in many scenarios, they still have some following limitations: (1) GraphRAG detects communities by relying on the Leiden algorithm, which however only uses the graph structure without considering the rich semantic information carried by the nodes and edges. As a result, the detected communities often consist of various different themes, which leads to the poor quality of community summaries and further decreases its performance. Besides, it uses LLMs to analyze all the communities when answering abstract questions, which is both time- and tokens-consuming. For example, the Leiden algorithm can detect 2,984 communities in the Multihop-RAG dataset, and to answer 100 questions, Global Search costs approximately \$650 and 106 million tokens [^1] The cost of GPT-4o is \$10/M tokens for output and \$2.50/M tokens for input (for details, please refer to openai.com/api/pricing)., which is very token-consuming. (2) LightRAG still cannot accurately identify the important keywords from the questions, which greatly affects its performance, and it does not generate community summaries in advance, making it difficult for the LLM to generate comprehensive summaries of all retrieved information. (3) HippoRAG fails to capture relationships between different chunks, making it susceptible to the “lost in the middle” dilemma (Liu et al., 2024b), which significantly prevents its ability to answer abstract questions, as it struggles to synthesize information across multiple chunks or understand overarching themes.
To tackle the above limitations, in this paper we propose a novel graph-based RAG approach, called Attributed Community-based Hierarchical RAG (ArchRAG), by augmenting the question using attributed communities from the knowledge graph built on the external corpus. Unlike GraphRAG which detects communities purely using links, we propose to detect communities by exploiting both links and the attributes of nodes, and call them attributed communities. Intuitively, an attributed community represents a group of nodes that are not only densely connected but also share similar themes, which provides detailed information for answering specific questions, and its summary offers a condensed view (Sarthi et al., 2024; Angelidis and Lapata, 2018), enabling more focused engagement with the content and making it suitable for answering abstract questions. To effectively answer both abstract and specific questions, we then organize the attributed communities with summaries into a hierarchical structure, which provides different granularity of abstractions for the whole knowledge graphs. Besides, we develop a novel hierarchical data structure for searching the nearest neighbors and the corresponding efficient online retrieval method.
Specifically, to detect attributed communities that are organized in a hierarchical manner, we propose a novel LLM-based iterative framework, which can incorporate any existing attributed community detection method (Zhou et al., 2009; Traag et al., 2019; Von Luxburg, 2007; Xu et al., 2007; Grover and Leskovec, 2016). In each iteration, we detect the attributed communities by considering both links and node attributes using a community detection method and generate a summary for each attributed community by using LLMs. Afterwards, we build a new smaller attributed graph by abstracting each attributed community as a single node and then linking each pair of nodes together if their summary similarity exceeds a threshold. By iterating the above steps multiple times, we can obtain a set of attributed communities that can be organized into a hierarchical tree structure.
Given the attributed communities organized hierarchically, to answer a question, a naive approach is to enumerate all the attributed communities and find the most relevant ones. However, this approach, in line with that of GraphRAG, is very time and tokens-consuming. To alleviate these issues, we propose to build a novel hierarchical index structure by borrowing some ideas from HNSW (Malkov and Yashunin, 2018), which is a well-known approximate nearest neighbor (ANN) search method, and we call it C-HNSW (Community-based HNSW). C-HNSW index facilitates the fast retrieval of the $k$ most relevant elements from each layer of the hierarchy, optimizing the RAG process. Additionally, by using adaptive filtering, we identify the most relevant information, ensuring that complementary data are retrieved across all layers. Through hierarchical search, we achieve cost-efficient generation and optimize the retrieval process without sacrificing the quality of the generated output.
We have extensively evaluated ArchRAG by conducting various experiments on real-world datasets, and the results show that ArchRAG outperforms existing approaches in answering both abstract and specific questions. Particularly, ArchRAG achieves a 10% high accuracy compared to state-of-the-art graph-based RAG methods when answering specific questions and also demonstrates significant improvement over others in answering abstract questions. Moreover, ArchRAG is very token-efficient, saving up to 250 times the token usage compared to GraphRAG (Edge et al., 2024).
In summary, our main contributions are as follows:
  • We present a novel graph-based RAG approach by using attributed communities that are organized hierarchically and detected by an LLM-based hierarchical clustering method.
  • To index attributed communities, we propose a novel hierarchical index structure called C-HNSW and also develop an efficient online retrieval method.
  • Extensive experiments show that ArchRAG is both highly effective and efficient, and achieves state-of-the-art performance in answering both abstract and specific QA tasks.
2. Preliminaries
Figure 2. ArchRAG consists of two phases: offline indexing and online retrieval. For the online retrieval phase, we show an example of using ArchRAG to answer a question in the HotpotQA dataset.
Existing graph-based RAG methods often follow a general framework that leverages external structured knowledge graphs to improve contextual understanding of LLMs and generate more informed responses (Peng et al., 2024). Typically, it involves two phases:
  1. (1)
    Offline indexing: Building a KG $G(V,E)$ from a given corpus $D$ where each vertex represents an entity and each edge denotes the relationship between two entities, and constructing an index based on the KG.
  2. (2)
    Online retrieval: Retrieving the relevant information (e.g., nodes, subgraphs, or textual information) from KG using the index and providing the retrieved information to the LLM to improve the accuracy of the response.
3. Our Approach ArchRAG
Unlike GraphRAG, which purely uses links to generate communities, we propose to detect communities that can be organized into a hierarchical tree structure by exploiting both links and attributes in the knowledge graph, also called attributed communities. An attributed community is a group of entities with a summary, where the entities are not only densely connected but also share similar themes, providing detailed information for answering specific questions, and the summary can offer a condensed view for answering abstract questions since it enables more focused engagement with the entities’ attributes. Moreover, in the hierarchical tree structure, the lower-level communities and entities contain detailed knowledge from the KG, while higher-level communities provide global summaries, which naturally enable ArchRAG to address questions with different granularity in abstraction.
Moreover, to answer questions, we need to efficiently identify highly relevant attributed communities at different levels of the hierarchical structure. Considering that our hierarchical attributed communities and the HNSW index (Malkov and Yashunin, 2018) are similar in structure, we propose to map entities and attributed communities into high-dimensional nodes and then build a unified Community-based HNSW (C-HNSW) index for them, which allows us to efficiently identify the highly relevant attributed communities.
As illustrated in Figure 2, our proposed ArchRAG consists of two phases. In the offline indexing phase, ArchRAG first constructs a KG from the corpus, then detects attributed communities by a novel LLM-based hierarchical clustering method, and finally builds the C-HNSW index. During the online retrieval phase, ArchRAG first converts the question into a query vector, then retrieves relevant information from the C-HNSW index, and finally generates answers through adaptive filtering-based generation process which identifies the most relevant information.
3.1. Offline Indexing
3.1.1. KG construction.
Our first step is to build a KG by prompting the LLM to extract entities and relations from each chunk of the text corpus $D$. Specifically, all text contexts are segmented into chunks based on specified chunk length, enabling the LLM to extract entities and relations from each chunk using in-context learning (Brown, 2020), thus forming subgraphs. These subgraphs are then merged, with entities and relations that appear repeatedly across multiple subgraphs being consolidated by the LLM to generate a complete description. Finally, we get a KG, denoted by $G(V,E)$, where $V$ and $E$ are sets of vertices and edges, respectively, and each vertex and edge is associated with some textual attributes.
3.1.2. LLM-based hierarchical clustering.
As aforementioned, a high-quality hierarchical attributed community structure provides different granularities of abstraction for the entire knowledge graph, which is beneficial for answering both abstract and specific questions. Therefore, we propose an iterative LLM-based hierarchical clustering framework that supports arbitrary graph augmentation (e.g., KNN connections and CODICIL (Ruan et al., 2013)) and graph clustering algorithms (e.g., weighted Leiden (Traag et al., 2019), weighted spectral clustering, and SCAN (Xu et al., 2007)). Specifically, we propose to augment the KG by linking entities if their attribute similarities are larger than a threshold (e.g., KNN connections), and then associate each pair of linked entities with a weight denoting their attribute similarity value. Next, we generate the attributed communities using any given graph clustering algorithm.
Algorithm 1 shows the above iterative clustering process. Given a graph augmentation method and clustering algorithm, we perform the following steps in each iteration: (1) augmenting the graph (line 3); (2) computing the edge weights (lines 4-5); (3) clustering the augmented graph (line 6); (4) generating a summary for each community using LLM (line 7); and (5) building a new attributed graph where each node denotes an attributed community and two nodes are linked if their community members are connected (line 9). We repeat the iterations until the stopping condition (such as insufficient nodes or reaching the specified level limit) is met. Since each iteration corresponds to one layer, all the attributed communities $HC$ can be organized into a multi-layer hierarchical tree structure, denoted by $\Delta$, where each community in one layer includes multiple communities in the next layer.
Appendix A.1 provides a running example for Algorithm 1.
input : KG ${G}(V,E)$, augment function Aug, graph clustering algorithm GCluster, and termination condition $T$
1 $T\leftarrow$ False, ${HC}\leftarrow\emptyset$;
2 repeat
3       ${G}^{\prime}(V,E^{\prime})\leftarrow$ Aug(${G}(V,E)$);
4
5      for  each $e^{\prime}=(u,v)\in E^{\prime}$ do
6             update the weight of $e^{\prime}$ as $1-\cos(z_{u},z_{v})$;
7
8      $C\leftarrow$ GCluster(${G}^{\prime}(V,E^{\prime})$);
9       for each $c\in C$ do  generate the summary of $c$ by LLM ;
10       ${HC}\leftarrow{HC}\cup C$;
11
12      ${G}(V,E)\leftarrow$ build a new graph using $C$ and $E^{\prime}$;
       // update $T$ according to $G(V,E)$;
13
14until $T$=True;
return ${HC}$;
Algorithm 1 LLM-based hierarchical clustering
3.1.3. C-HNSW index.
Given a query, to efficiently identify the most relevant information from each layer of the hierarchical tree $\Delta$, a naive method is to build a vector database for the attributed communities in each layer, which is costly in both time and space. To tackle this issue, we propose to build a single hierarchical index for all the communities. Recall that the attributed communities in $\Delta$ form a tree structure, and the number of nodes decreases as the layer level increases. Since this tree structure is similar to the HNSW (Hierarchical Navigable Small World) index which is the most well-known index for efficient ANN search (Malkov and Yashunin, 2018) (Appendix A.2 introduces its details), we propose to map entities and attributed communities of $\Delta$ into high-dimensional nodes, and then build a unified Community-based HNSW (C-HNSW) index for them.
$\bullet$ The structure of C-HNSW. Conceptually, the C-HNSW index is a list of simple graphs with links between them, denoted by $\mathcal{H}=({\mathcal{G}},{L_{inter}})$ with $\mathcal{\mathcal{G}}=\{G_{0}=(V_{0},E_{0}),G_{1}=(V_{1},E_{1}),\cdots,G_{L}=(V_{L},E_{L})\}$, where $G_{i}$ is a simple graph and each node of the simple graph corresponds to an attributed community or entity. The number $L$ of layers (simple graphs) of $\mathcal{H}$ is the same as the that of $\Delta$. Specifically, for each attributed community or entity in the $i$-the layer of $\Delta$, we map it to a high-dimensional node in the $i$-th layer of $\mathcal{H}$ by using a language model (e.g., nomic-embed-text (Nussbaum et al., 2024)).
After obtaining the high-dimensional nodes, we establish two types of links between them, i.e., intra-layer and inter-layer links:
  • Intra-layer links: These links exist between nodes in the same layers. Specifically, for each node in each layer, we link it to at least $M$ nearest neighbors within the same layer, where $M$ is a predefined value, and the nearest neighbors are determined according to a given distance metric $d$. Thus, all the intra-layer links are edges in all the simple graphs:
    (1) $L_{intra}=\bigcup_{i=0}^{L}E_{i}$
  • Inter-layer links: These links cross two adjacent layers. Specifically, we link each node in each layer to its nearest neighbor in the next layer. As a result, all the inter-layer links can be represented as follows:
    (2) $L_{inter}=\bigcup_{i=1}^{L}\{(v,\psi(v))|v\in V_{i},\psi(v)\in V_{i-1}\},$
    where $\psi(\cdot):V_{i}\rightarrow V_{i-1}$ is the injective function that identifies the nearest neighbor of each node in the lower layer.
For example, in Figure 2, the C-HNSW index has three layers (simple graphs), incorporating all the attributed communities. Within each layer, each node is connected to its two nearest neighbors via intra-layer links, denoted by undirected edges. The inter-layer links are represented by arrows, e.g., the green community at layer $L_{1}$ is connected to the green entity (its nearest neighbor at layer $L_{0}$).
Intuitively, since the two types of links above are established based on nearest neighbors, C-HNSW allows us to quickly search the relevant information for a query by traversing along with these links. Note that C-HNSW is different from HNSW since it has intra-layer links and each node exists in only one layer.
$\bullet$ The construction of C-HNSW. A naive approach to build the C-HNSW index is to build nodes first and then establish the two types of links by finding the nearest neighbors of each node. However, the process of finding the nearest neighbor is costly. To accelerate the construction, we propose a top-down approach by borrowing the idea of HNSW construction. Specifically, by leveraging the query process of C-HNSW, which will be introduced in Section 3.2.1, nodes are progressively inserted into the index starting from the top layer, connecting intra-layer links within the same layer and updating the inter-layer links. For lack of space, we give the details of the construction algorithm in Appendix A.2.
3.2. Online retrieval
In the online retrieval phase, after obtaining the query vector for a given question, ArchRAG generates the final answer by first conducting hierarchical search on the C-HNSW index, and then analyzing and filtering of retrieved information.
3.2.1. Hierarchical search.
We first introduce the query process of C-HNSW and then describe the hierarchical search. The query process is the core of C-HNSW, which can be used for constructing C-HNSW as illustrated in Algorithm 2.
input : $\mathcal{H}=({\mathcal{G}},{L_{inter}})$, $q$, $k$, $l$.
1 $s\leftarrow$ a random node in the highest layer $L$;
2 for $i\leftarrow L,\cdots,l+1$ do
3       $c\leftarrow$ SearchLayer(${G_{i}}=(V_{i},E_{i}),q,s,1,i$);
4       $s\leftarrow$ find the node in layer $i-1$ via the inter-layer links of $c$;
5
[…]
7 return $K$;
8Procedure SearchLayer(${G_{i}}=(V_{i},E_{i}),q,s,k,i$);
9 $V\leftarrow\{s\}$, $R\leftarrow\{s\}$, $Q\leftarrow$ initialize a queue containing $s$;
10 while $|R|>0$ do
11       $c\leftarrow$ nearest node in $Q$;
12       $f\leftarrow$ furthest node in $R$;
13       if $d(c,q)>d(f,q)$ then  break ;
14       for each neighbor $x\in N(c)$ in $G_{i}$ do
15             if $x\in V$ then continue;
16             $V\leftarrow V\cup\{x\}$;
17             $f\leftarrow$ furthest node in $R$;
18             if $d(x,q)<d(f,q)$ or $|R|<k$  then
19                   $Q\leftarrow Q\cup\{x\}$, $R\leftarrow R\cup\{x\}$;
20                   if $|R|>k$ then remove $f$ from $R$;
21
return ${R}$;
Algorithm 2 C-HNSW query
Specifically, given a specified query layer $l$, query point $q$, and the number $k$ of nearest neighbors, the query algorithm can be implemented through the following iterative process:
  1. (1)
    Start from a random node at the highest layer $L$, which serves as the starting node for layer $L$ (line 1).
  2. (2)
    For each layer from layers $L$ to $l+1$, begin at the starting node and use a greedy traversal approach (i.e., the procedure SearchLayer) to find the nearest neighbor $c$ of $q$, and then traverse to the next layer using $c$’s inter-layer link as the starting node for the next layer (lines 2-4).
  3. (3)
    In the query layer $l$, use the greedy traversal approach to find the $k$ nearest neighbors of $q$ (line 5).
Specifically, the greedy traversal strategy compares the distance between the query point and the visited nodes during the search process. It achieves this by maintaining a candidate expansion queue $Q$ and a dynamic nearest neighbor set $R$, which contains $k$ elements:
  • Expansion Queue $Q$: For each neighbor $x$ of a visited node, if $d(x,q)<d(f,q)$, where $f$ is the furthest node from $R$ to $q$, then $x$ is added to the expansion queue.
  • Dynamic Nearest Neighbor Set $R$: Nodes added to $C$ are used to update $R$, ensuring that it maintains no more than $k$ elements, where $k$ is the specified number of query results.
In the greedy traversal, if a node $x$ expanded from $Q$ satisfies $d(n,q)>d(n,f)$, where $f$ is the furthest node from $R$ to $q$, then the traversal stops.
Based on the discussions above, we propose a hierarchical search approach, which retrieves the top $k$ relevant elements in each layer using the method outlined in Algorithm 2. Once the query at a given layer is complete, we use the inter-layer link of the nearest neighbor in that layer to traverse to the next layer, repeating the process until reaching the bottom layer. At the bottom layer, we extract the relationships between connected retrieved entities, forming the text information $R_{0}$ for this subgraph. We represent the retrieved information from each layer as $R_{i}$, where $i\in\{0,1,\cdots,L\}$, to be used in the adaptive filtering-based generation process.
3.2.2. Adaptive filtering-based generation.
While some optimized LLMs support longer text inputs, they may still encounter issues such as the “lost in the middle” dilemma (Liu et al., 2024b). Thus, direct utilization of retrieved information comprising multiple text segments for LLM-based answer generation risks compromising output accuracy.
To mitigate this limitation, we propose an adaptive filtering-based method that harnesses the LLM’s inherent reasoning capabilities. We first prompt the LLM to extract and generate an analysis report from the retrieved information, identifying the parts that are most relevant to answering the query and assigning relevance scores to these reports. Then, all analysis reports are integrated and sorted, ensuring that the most relevant content is used to summarize the final response to the query, with any content exceeding the text limit being truncated. This process can be represented as:
(3) $A_{i}$ $=LLM(P_{filter}||R_{i})$
(4) $Output$ $=LLM(P_{merge}||Sort(\{A_{0},A_{1},\cdots,A_{n}\}))$
where $P_{filter}$ and $P_{merge}$ represent the prompts for extracting relevant information and summarizing, respectively, $A_{i}$, $i\in 0\cdots n$ denotes the filtered analysis report. The sort function orders the content based on the relevance scores from the analysis report.
We also analyze the complexity of ArchRAG in Appendix A.3.
4. Experiments
Our experiments aim to answer following research questions (RQs):
  • RQ1: How does ArchRAG perform compared with existing baselines on specific and abstract QA tasks?
  • RQ2: How efficient is the ArchRAG approach?
  • RQ3: What is the quality of our LLM-based attributed communities, and how does it impact the performance of RAG?
  • RQ4: How robust is ArchRAG?
4.1. Setup
Table 2. Datasets used in our experiments. BLEU, MET, and ROU denote BLEU-1, METEOT, and ROUGE-L F1.
Dataset Multihop-RAG HotpotQA NarrativeQA
Passages 609 9,221 1,572
Tokens 1,426,396 1,284,956 121,152,448
Nodes 23,353 37,436 650,571
Edges 30,716 30,758 679,426
Questions 2,556 1,000 43,304
Metrics Accuracy, Recall Accuracy, Recall BLEU, MET, ROU
Datasets.
We consider both specific and abstract QA tasks. For the specific QA task, we use the following datasets: Multihop-RAG (Tang and Yang, 2024), HotpotQA (Yang et al., 2018), and NarrativeQA (Kočiskỳ et al., 2018), all of which are extensively utilized within the QA and Graph-based RAG research communities (Hu et al., 2022; Gutiérrez et al., 2024; Yao et al., 2022; Xu et al., 2024; Sarthi et al., 2024; Asai et al., 2023). Multihop-RAG is designed to evaluate retrieval and reasoning across news articles in various categories. HotpotQA is a QA dataset featuring natural, multi-hop questions. NarrativeQA is a dataset that comprises QA pairs based on the full texts of books and movie transcripts. Besides, we follow the GraphRAG (Edge et al., 2024) method for the abstract QA task. We reuse the Multihop-RAG corpus and prompt LLM to generate questions that convey a high-level understanding of dataset contents. The statistics of these datasets are reported in Table 2.
Table 3. Performance comparison of different methods across various datasets for solving specific QA tasks. The best and second-best results are marked in bold and underlined, respectively.
Baseline Type Method Multihop-RAG HotpotQA NarrativeQA
(Accuracy) (Recall) (Accuracy) (Recall) (BLEU-1) (METEOR) (ROUGE-L F1)
Inference-only Zero-shot 47.7 23.6 28.0 31.8 8.0 7.9 8.6
CoT 54.5 28.7 32.5 39.7 5.0 8.1 6.4
Retrieval-only BM25 37.6 19.4 49.7 53.6 2.0 4.9 2.8
Vanilla RAG 58.6 31.4 50.6 56.1 2.0 4.9 2.8
Graph-based RAG HippoRAG 38.9 19.1 51.3 56.8 2.2 5.0 2.8
LightRAG-Low 44.1 25.1 34.1 41.8 4.5 8.7 6.6
LightRAG-High 48.5 28.7 25.6 33.3 4.4 8.1 6.1
LightRAG-Hybrid 50.3 30.3 35.6 43.3 5.0 9.4 7.0
GraphRAG-Local 40.1 23.8 29.7 35.5 3.9 3.3 3.5
GraphRAG-Global 45.9 28.4 33.5 42.6 2.9 3.7 3.3
Our proposed ArchRAG 68.8 37.2 65.4 69.2 11.5 15.6 17.6
Baselines.
Our experiments consider three model configurations:
  • Inference-only: Using an LLM for direct question answering without any retrieval data, i.e., Zero-Shot and CoT (Kojima et al., 2022).
  • Retrieval-only: Retrieval models extract relevant chunks from all documents and use them as prompts for large models. We select strong and widely used retrieval models: BM25 (Robertson and Walker, 1994) and vanilla RAG.
  • Graph-based RAG: These methods leverage graph data during retrieval. We select HippoRAG (Gutiérrez et al., 2024), GraphRAG (Edge et al., 2024), and LightRAG (Guo et al., 2024). Particularly, GraphRAG has two versions, i.e., GraphRAG-Global and GraphRAG-Local, which use global and local search methods respectively. Similarly, LightRAG integrates local search, global search, and hybrid search, denoted by LightRAG-Low, LightRAG-High, and LightRAG-Hybrid, respectively.
In GraphRAG-Global, all communities below the selected level are first retrieved, and then the LLM is used to filter out irrelevant communities. This process can be viewed as utilizing the LLM as a retriever to find relevant communities within the corpus. According to the selected level of communities (Edge et al., 2024), GraphRAG-Global can be further categorized into C1 and C2, representing high-level and intermediate-level communities, respectively, with C2 as the default.
Metrics.
For the specific QA tasks, we use Accuracy and Recall to evaluate performance on the first two datasets based on whether gold answers are included in the generations instead of strictly requiring exact matching, following (Schick et al., 2024; Mallen et al., 2022; Asai et al., 2023). We also use the official metrics of BLEU, METEOR, and ROUGE-l F1 in the NarrativeQA dataset. For the abstract QA task, we follow prior work (Edge et al., 2024) and adopt a head-to-head comparison approach using an LLM evaluator (GPT-4o). LLMs have demonstrated strong capabilities as evaluators of natural language generation, often achieving state-of-the-art or competitive results when compared to human judgments (Wang et al., 2023b; Zheng et al., 2023). Overall, we utilize four evaluation dimensions: Comprehensiveness, Diversity, Empowerment, and Overall. Please see Appendix B.1 for more details on metrics.
Implementation details.
We mainly use Llama 3.1-8B (Dubey et al., 2024) as the default LLM model and use nomic-embed-text (Nussbaum et al., 2024) as the embedding model to encode all texts. We use KNN for graph augmentation and the weighted Leiden algorithm for community detection. For each retrieval item $k$, we search the same number of items at each layer, with $k=5$ as the default. Our codes are available at: https://anonymous.4open.science/r/H-CAR-AG-1E0B. Additional details are provided in Appendix B.2.
4.2. Overall results
To answer the question RQ1, we compare our method with baseline methods in solving both specific and abstract QA tasks.
$\bullet$ Results of specific QA tasks. Table 3 reports the performance of each method on three datasets. Clearly, ArchRAG demonstrates a substantial performance advantage over other baseline methods on these datasets. On the Multihop-RAG dataset, HippoRAG performs worse than retrieval-only methods while outperforming retrieval-based methods on the HotpotQA dataset. This is mainly because, on the HotpotQA dataset, passages are segmented by the expert annotators; that is, passages can provide more concise information, whereas on Multihop-RAG, passages are segmented based on chunk size, which may cause the LLM to lose context and produce incorrect answers. The experimental results suggest that not all communities are suitable for specific QA tasks, as the GraphRAG-Global performs poorly. Furthermore, GraphRAG does not consider node attributes during clustering, which causes the community’s summary to become dispersed, making it difficult for the LLM to extract relevant information. Thus, we gain an interesting insight: LLM may not be a good retriever, but is a good analyzer.
$\bullet$ Results of abstract QA tasks. We compare ArchRAG against baselines across four dimensions on the Multihop-RAG dataset. Here, we only compare the LightRAG-Hybrid method, as it represents the best version of LightRAG methods (Guo et al., 2024). As shown in Figure LABEL:fig:winmap, our method achieves comparable performance with GraphRAG-Global on the diversity and empowerment dimensions while significantly surpassing it on the comprehensive dimension. Overall, ArchRAG demonstrates outstanding performance in solving abstract QA tasks.
$\bullet$ Efficiency of ArchRAG. To answer the question RQ2, we compare the time cost and token usage of ArchRAG with those of other baseline methods and the index consumption against that of graph-based RAG methods. As shown in Figure LABEL:fig:query-efficiency, ArchRAG demonstrates significant time and cost efficiency for online queries. For example, on the HotpotQA dataset, ArchRAG reduces token usage by a factor of 250 compared to GraphRAG-Global, utilizing 5.1M tokens and 1,394M tokens, respectively.
LABEL:eff_leg
Besides, Figure LABEL:fig:index-efficiency shows the index construction time and token usage. The cost of building index for ArchRAG is similar to that of GraphRAG, but due to the need for community summarization, both take higher time cost and token usage than HippoRAG.
LABEL:eff_index
4.3. Detailed Analysis
To answer the question RQ3, we first evaluate the community qualities under our proposed LLM-based attributed clustering framework and then conduct an ablation study.
$\bullet$ Community quality of different clustering methods. We evaluate 4 different clustering methods with 2 graph augmentation methods. The clustering methods we choose are weighted Leiden, Spectral Clustering (Von Luxburg, 2007), SCAN (Xu et al., 2007), and node2vec (Grover and Leskovec, 2016) with KMeans. The graph augmentation methods we choose are the KNN algorithm, which computes the similarity between each node, and CODICIL (Ruan et al., 2013), which selects and adds the top similar edges to generate better clustering results. We utilize the Calinski-Harabasz (CH) Index (Caliński and Harabasz, 1974) and Cosine Similarity (Charikar, 2002), abbreviated as CHI and Sim respectively, to evaluate the quality of generated attributed communities. The higher CH Index and Cosine Similarity indicate a higher quality of the community. The details of evaluation metrics can be found in Appendix B.1. As shown in Figures LABEL:fig:ch and LABEL:fig:sim, both KNN and CODICIL help generate high-quality communities well.
LABEL:quality_ch
LABEL:quality_sim
$\bullet$ Ablation study. To evaluate the effect of the attributed community in ArchRAG, we design some variants of ArchRAG and conduct the ablation experiments.
  • Spec.: Replaces the Weighted Leiden community detection algorithm with spectral clustering.
  • Spec. (No Aug): Removes the graph augmentation and uses spectral clustering.
  • Leiden: Replaces our clustering framework with the Leiden algorithm.
  • Single-Layer: Replaces our hierarchical index and search with a single-layer community.
  • Entity-Only: Generate the response using entities only.
  • Direct Prompt: Direct prompts the LLM to generate the response without the adaptive filtering-based generation.
The first two variants replace the algorithm in the LLM-based hierarchical clustering framework, while the last three variants modify key components of ArchRAG: attributes, hierarchy, and communities. As shown in Table 4, the performance of ArchRAG on specific QA tasks decreases when each feature is removed, with the removal of the community component resulting in the most significant drop. Additionally, the Direct variant demonstrates that the adaptive filtering-based generation process can effectively extract relevant information from retrieved elements.
Table 4. Comparing the performance of different variants of ArchRAG on the specific QA tasks.
Method variants Multihop-RAG HotpotQA
(Accuracy) (Recall) (Accuracy) (Recall)
ArchRAG 68.8 37.2 65.4 69.2
- Spec. 67.1 36.7 60.5 63.7
- Spec. (No Aug) 62.8 34.2 64.8 63.2
- Leiden 63.2 34.1 61.7 64.8
- Single-Layer 63.8 36.4 60.1 63.6
- Entity-Only 61.2 34.7 59.9 63.1
- Direct Prompt 59.9 29.6 40.7 45.4
In addition, we evaluate the performance of ArchRAG under different LLM backbones and various top-k retrieval content to address question RQ4.
$\bullet$ Effect of LLM backbones. As strong LLMs with hundreds of billions of parameters (e.g., GPT-3.5-turbo) possess enhanced capabilities, our proposed ArchRAG may also benefit from performance improvement. As shown in Table 5, the results of Llama 3.1-8B are similar to those of GPT-3.5-turbo, as Llama3.1’s capabilities are comparable to those of GPT-3.5-turbo (Dubey et al., 2024). Besides, GPT-4o-mini performs better than other LLM backbones because of its exceptional reasoning capabilities.
Table 5. Comparing ArchRAG with other RAG methods on the specific QA tasks under different LLM backbone models.
LLM backbone Multihop-RAG HotpotQA
(Accuracy) (Recall) (Accuracy) (Recall)
Llama3.1-8B 68.8 37.2 65.4 69.2
GPT-3.5-turbo 67.2 31.5 62.8 65.0
GPT-4o-mini 77.3 33.8 69.9 73.8
$\bullet$ Effect of $k$ values. We compare the performance of ArchRAG under different retrieved elements. As shown in Figure LABEL:fig:topk, the performance of ArchRAG shows little variation when selecting different retrieval elements (i.e., communities and entities in each layer). This suggests that the adaptive filtering process can reliably extract the most relevant information from the retrieval elements and integrate it to generate the answer.
LABEL:topk_leg
5. Related Work
In this section, we review the related works, including Retrieval-Augmentation-Generation (RAG) approaches, and LLMs for graph mining and learning.
$\bullet$ RAG approaches. RAG has been proven to excel in many tasks, including open-ended question answering (Jeong et al., 2024; Siriwardhana et al., 2023), programming context (Chen et al., 2024a, 2023), SQL rewrite (Li et al., 2025; Sun et al., 2024), and data cleaning (Naeem et al., 2024; Narayan et al., 2022; Qian et al., 2024). The naive RAG technique relies on retrieving query-relevant contexts from external knowledge bases to mitigate the “hallucination” of LLMs. Recently, many RAG approaches (Edge et al., 2024; Guo et al., 2024; Gutiérrez et al., 2024; Sarthi et al., 2024; Wang et al., 2024a; He et al., 2024; Ma et al., 2024, 2024; Li et al., 2024) have adopted graph structures to organize the information and relationships within documents, achieving improved overall retrieval performance. For more details, please refer to the recent survey of graph-based RAG methods (Peng et al., 2024).
$\bullet$ LLM for graph mining. Recent advances in LLMs have offered opportunities to leverage LLMs in graph mining. These include using LLMs for KG construction (Zhu et al., 2024), and employing KG to enhance the LLM reasoning (Wang et al., 2023a, 2024a; Luo et al., 2023; Sun et al., 2023; Mavromatis and Karypis, 2024; Jiang et al., 2023). For instance, KD-CoT (Wang et al., 2023a) retrieves facts from an external knowledge graph to guide the CoT performed by LLMs. RoG (Luo et al., 2023) proposes a planning-retrieval-reasoning framework that retrieves reasoning paths from KGs to guide LLMs conducting faithful reasoning. To capture graph structure, GNNRAG (Mavromatis and Karypis, 2024) adopts a lightweight graph neural network to retrieve elements from KGs effectively. StructGPT (Jiang et al., 2023) and ToG (Sun et al., 2023) treat LLMs as agents that interact with KGs to find reasoning paths leading to the correct answers. Besides, LLM is used to generate code to solve complex graph mining problems, e.g., the shortest path computation (Chen et al., 2024b; Tang et al., 2024; Cao et al., 2024; Zhang et al., 2024a).
$\bullet$ LLM for graph learning. Recently, LLMs have demonstrated remarkable zero-shot capabilities in processing text-attributed graphs. Following this trend, graph foundation models have emerged as a promising direction (Liu et al., 2024a; Zhang et al., 2024b). For instance, Liu et al. (Liu et al., 2024a) proposed a unified model capable of handling various graph ML tasks, such as node classification and link prediction, thereby eliminating the need for task-specific architectures.
6. Conclusion
In this paper, we propose ArchRAG, a novel graph-based RAG approach, by augmenting the question using attributed communities from the knowledge graph built on the external corpus, and building a novel index for efficient retrieval of relevant information. Our experiments show that ArchRAG is highly effective and efficient. In the future, we will explore fast parallel graph-based RAG methods to process large-scale external corpus.
Appendix A Method details
A.1. LLM-based hierarchical clustering
For example, as shown in the second step of offline indexing in Figure 2, LLM-based hierarchical clustering first enhances the original KG at layer $L_{0}$ by adding similar edges. Next, the weight of each edge is calculated based on the strength of the relationship between the node embeddings, and a weighted clustering algorithm is applied to obtain four communities. Based on the existence of links between nodes within a community, the topology of communities at layer $L_{1}$ is constructed, resulting in a graph of communities. This process is repeated, ultimately generating a clustering result consisting of three layers of hierarchical communities.
A.2. More details of C-HNSW
$\bullet$ Introduction of HNSW. We provide a brief introduction to HNSW, an efficient Approximate Nearest Neighbor Search (ANNS) technique for vector databases.
Definition 1 (Hierarchical Navigable Small World (HNSW) (Malkov and Yashunin, 2018)).
HNSW is a graph-based ANNS algorithm that consists of a multi-layered index structure, where each node uniquely corresponds to a vector in the database. Given a set $S$ containing $n$ vectors, the constructed HNSW can be represented as a pair $\mathcal{H}=(\mathcal{\bf G},\mathcal{C})$. $\mathcal{\bf G}=\{G_{0},G_{1},\dots,G_{L}\}$ is a set of simple graphs (also called layers) $G_{i}=(V_{i},E_{i})$, where $i\in\{0,1,\dots,L\}$ and $V_{L}\subset V_{L-1}\subset\dots\subset V_{1}\subset V_{0}=S$. $\mathcal{C}$ records the inter-layer mappings of edges between the same node across adjacent layers $\mathcal{C}=\bigcup_{i=0}^{L-1}\{(v,\phi(v))|v\in V_{i},\phi(v)\in V_{i+1}\}$, where $\phi(v):V_{i}\rightarrow V_{i+1}$ is the mapping function for the same node across two adjacent layers.
The nodes in the multi-layer graph of HNSW are organized in a nested structure, where each node at each layer is connected to its nearest neighbors. During a query, the search begins at the top layer and quickly identifies the node closest to $q$ through a greedy search. Then, through inter-layer mapping, the search proceeds to the next lower layer. This process continues until all approximate nearest neighbors are identified in $G_{0}$.
$\bullet$ The construction of C-HNSW. The construction of C-HNSW is illustrated in Algorithm 3. The construction algorithm of C-HNSW follows a top-down approach. Using the query process of C-HNSW, we obtain the $M$ nearest neighbors of each node in its layer, and the inter-layer links are continuously updated during this process. When inserting a node $x$ at layer $i$ , if the nearest neighbor at the layer $i+1$ is $c_{j}$, the inter-layer link of $c_{j}$ is updated in the following two cases:
  • Node $c_{j}$ does not have a inter-layer link to the layer $i$.
  • The distance from node $c_{j}$ to node $x$ is smaller than the distance from $c_{j}$ to its previous nearest neighbor $x^{\prime}$, i.e., $d(c_{j},x)<d(c_{j},x^{\prime})$.
After all nodes at layer $i$ ($i<L$) have been inserted, we check each node at layer $i+1$ to ensure it has a inter-layer link, confirming the traverse from the higher layer to the lower layer.
input : The Hierarchical community $\mathcal{HC}$, KG $\mathcal{G}(V,E)$, maximum number of connections for each node $M$.
1 $\mathcal{H}\leftarrow\emptyset$;
2
${\bf V}\leftarrow$$\mathcal{HC}\cup V$} // Get all nodes of each layer.
3 for each layer $l\leftarrow L\cdots 0$ do
4       for each node $v\in V_{l}$ do
5             if $l\neq L$ then
6                   $R\leftarrow$ SearchLayer(${G_{l}}=(V_{l},E_{l}),q,s,1,l$);
7                   $c\leftarrow$ get the nearest node from R;
8                   $s\leftarrow$ node in layer $l-1$ via $c$’s inter-layer link;
9                   if $s$ is null or $d(c,s)>d(c,v)$ then
10                         update $v$ as $c$’s inter-layer link
11            if $l=L$ or $s$ is null then $s\leftarrow$ random node in layer $l$;
12             $R\leftarrow$ SearchLayer(${G_{l}}=(V_{l},E_{l}),q,s,M,l$);
13             add edges between $v$ and $R$, update $E_{l}$;
14
15      $\mathcal{H}\leftarrow\mathcal{H}\cup{G_{l}}=(V_{l},E_{l})$
return $\mathcal{H}$;
Algorithm 3 C-HNSW construction
A.3. Complexity analysis of ArchRAG
We now analyze the complexity of our ArchRAG approach. Since the token cost is more important in the era of LLM, we replace the space complexity with the token cost.
The offline indexing process of ArchRAG includes the KG construction, hierarchical clustering, and C-HNSW construction. The time complexity and token usage are as follows:
Lemma 2.
Given a large text corpus or a large set of text documents with a total of $D$ tokens, the time complexity of the offline indexing process of ArchRAG is $O(I\frac{D}{w}+\frac{1-a^{L}}{1-a}(n*t+I\frac{D}{w}+\pi(m)+n\log n))$, where ${\bf I}$ is the generation time of the LLM for a single inference, $w$ is the specified token size of one chunk, $n$ and $m$ are the number of entities and relations in the extracted KG, $L$ is the height of the resulting hierarchical community structure, and $a$ is the average ratio of the number of nodes between two consecutive layers, with $10<a<$ 1. For a given embedding model, the computation time for the embedding of an entity description is denoted by $t$, while a specific clustering method is typically a function of $m$, represented as $\pi(m)$.
Proof.
For the corpus $D$, we use the LLM to infer and extract the KG from each chunk of size $w$, resulting in a cost of $O(I\frac{D}{w})$ for constructing the KG. Since the size of all community summaries in a single layer would not exceed the length of the corpus, their LLM inference time is also less than $O(I\frac{D}{w})$. For each layer of clustering, the embedding of each point must be computed, and clustering is performed with time complexity of $\pi(m)$. In the C-HNSW construction, each point requires $O(nlogn)$ time to perform the k-nearest neighbor search and establish connections, which is similar to the proof in (Malkov and Yashunin, 2018). For an L-layer multi-layer graph structure, where the number of nodes decreases by a factor of $a$ between two consecutive layers, the increase in clustering and C-HNSW construction time is given by: $\frac{1-a^{L}}{1-a}$. ∎
Lemma 3.
Given a large text corpus or a large set of text documents with a total of $D$ tokens, the number of tokens used in the offline indexing process of ArchRAG is $O(D+D\frac{1-a^{L}}{1-a})$, where $L$ is the height of the resulting hierarchical community structure and $a$ is the average ratio of the number of nodes between two consecutive layers, with $10<a<$ 1.
Proof.
Based on the above proof, the token cost for constructing the KG is $O(D)$, while the token cost for the summaries does not exceed $O(D\frac{1-a^{L}}{1-a})$. ∎
Generally, $L$ is $O(\log n)$, where $n$ is the number of extracted entities. However, due to the constraints of community clustering, $L$ is typically constant, usually no greater than 5.
Next, we analyze the time complexity and token usage in the online query process of the ArchRAG.
Lemma 4.
Given a C-HNSW with $L$ layers constructed from a large text corpus or a large set of text documents, the time complexity of a sequentially executed single online retrieval query in ArchRAG is $O(e+LkI+Lk\log(n))$, where $I$ is the generation time of the LLM for a single inference, $e$ is the time cost of computing the query embedding, $k$ is the number of nodes retrieved at each layer, and $n$ is the number of nodes at the lowest layer in C-HNSW.
Proof.
ArchRAG first computes the embedding of the query, which takes $O(e)$ time. For each layer, querying one nearest neighbor takes no more than $O(\log(n))$ time, similar to the proof in (Malkov and Yashunin, 2018). In the Adaptive filtering-based generation, the content of each query is analyzed and inferred, requiring $O(LkI)$ time. Therefore, the total time for the online retrieval query is $O(e+LkI+Lk\log(n))$. ∎
Lemma 5.
Given a C-HNSW with $L$ layers constructed from a large text corpus or a large set of text documents, the number of tokens used for a single online retrieval query in ArchRAG is $O(kL(c+P)))$, where $k$ is the number of neighbors retrieved at each layer, $c$ is the average length of the retrieved content, and $P$ is the length of the prompt.
Proof.
The token consumption for analyzing all retrieved information is $O(kL(c+P)))$, while the token consumption for generating the final response is of constant order. Therefore, the total token consumption for the online retrieval is $O(kL(c+P)))$. ∎
In practice, multiple retrieved contents are combined, allowing the LLM to analyze them together, provided the total token count does not exceed the token size limit. As a result, the time and token usage for online retrieval are lower than those required for analysis.
Appendix B Experimental details
B.1. Metrics
This section provides additional details on the metrics.
$\bullet$ Metrics for specific QA tasks. We choose accuracy as the evaluation metric based on whether the gold answers are included in the model’s generations rather than strictly requiring an exact match, following (Schick et al., 2024; Mallen et al., 2022; Asai et al., 2023). This is because LLM outputs are typically uncontrollable, making it difficult for them to match the exact wording of standard answers. Similarly, we choose recall as the metric instead of precision, as it better reflects the accuracy of the generated responses. Additionally, when calculating recall, we adopt the same approach as previous methods (Gutiérrez et al., 2024; Asai et al., 2023): if the golden answer or the generated output contains “yes” or “no”, the recall for that question is set to 0. Therefore, the recall metric is not perfectly correlated with accuracy.
$\bullet$ Metrics for abstract QA tasks. Following existing works, we use an LLM to generate abstract questions, with the prompts shown in Figure 8, defining ground truth for abstract questions, particularly those involving complex high-level semantics, poses significant challenges. We build on existing works (Edge et al., 2024; Guo et al., 2024) to address this and adopt an LLM-based multi-dimensional comparison method (including comprehensiveness, diversity, empowerment, and overall). We employ a robust LLM, specifically GPT-4o-mini, to rank each baseline against our method. Figure 9 shows the evaluation prompt we use.
Prompt for generating abstract questions Prompt: Given the following description of a dataset: {description} Please identify 5 potential users who would engage with this dataset. For each user, list 5 tasks they would perform with this dataset. Then, for each (user, task) combination, generate 5 questions that require a high-level understanding of the entire dataset. Output the results in the following structure: - User 1: [user description]: : - Task 1: [task description]: : : - Question 1: : - Question 2: : - Question 3: : - Question 4: : - Question 5: - Task 2: [task description]: : - Task 5: [task description]: - User 2: [user description]: : - User 5: [user description]: : Note that there are 5 users and 5 tasks for each user, resulting in 25 tasks in total. Each task should have 5 questions, resulting in 125 questions in total. The Output should present the whole tasks and questions for each user. Output:
Figure 8. The prompt for generating abstract questions.
Prompt for LLM-based multi-dimensional comparison Prompt: You will evaluate two answers to the same question based on three criteria: Comprehensiveness, Diversity, Empowerment, and Directness. Comprehensiveness: How much detail does the answer provide to cover all aspects and details of the question? Diversity: How varied and rich is the answer in providing different perspectives and insights on the question? Empowerment: How well does the answer help the reader understand and make informed judgments about the topic? Directness: How specifically and clearly does the answer address the question? For each criterion, choose the better answer (either Answer 1 or Answer 2) and explain why. Then, select an overall winner based on these four categories. Here is the question:
    Question: {query}
Here are the two answers:
    Answer 1: {answer1}
    Answer 2: {answer2}
Evaluate both answers using the four criteria listed above and provide detailed explanations for each criterion. Output your evaluation in the following JSON format:
    "Comprehensiveness": {
        "Winner": "[Answer 1 or Answer 2]",
        "Explanation": "[Provide one sentence explanation here]"
    "Diversity": {
        "Winner": "[Answer 1 or Answer 2]",
        "Explanation": "[Provide one sentence explanation here]"
    "Empowerment": {
        "Winner": "[Answer 1 or Answer 2]",
        "Explanation": "[Provide one sentence explanation here]"
    "Overall Winner": {
        "Winner": "[Answer 1 or Answer 2]",
        "Explanation": "[Briefly summarize why this answer is the overall winner]"
Output:
Figure 9. The prompt for the evaluation of abstract QA.
$\bullet$ Metrics of community quality. We select the following metrics to evaluate the quality of the community:
  1. (1)
    Calinski-Harabasz Index (CHI) (Caliński and Harabasz, 1974): A higher value of CHI indicates better clustering results because it means that the data points are more spread out between clusters than they are within clusters. It is an internal evaluation metric where the assessment of the clustering quality is based solely on the dataset and the clustering results and not on external ground-truth labels. The CHI is calculated by between-cluster separation and within-cluster dispersion:
    (5) $CHI=\frac{N-C}{C-1}\frac{\sum^{C}_{i=1}n_{i}||c_{i}-c||^{2}}{\sum^{C}_{i=1}\sum_{x\in C_{i}}||x-c_{i}||^{2}}.$
    $N$ is the number of nodes. $C$ is the number of clusters.$n_{i}$ is the number of nodes in cluster $i$. $C_{i}$ is the $i-th$ cluster. $c_{i}$ is the centroid of cluster $C_{i}$. $c$ is the overall centroid of the datasets. $x$ is the feature of the target node.
  2. (2)
    Cosine Similarity (Sim) (Charikar, 2002): Cosine similarity is a measure of similarity between two non-zero vectors defined in an inner product space. In this paper, for each cluster, we calculate the similarity between the centroid of this cluster and each node in this cluster:
    (6) $Sim=\frac{1}{N}\sum^{C}_{i=1}\sum_{x\in C_{i}}Cosine(x,c_{i}).$
    $N$ is the number of nodes. $C$ is the number of clusters. $C_{i}$ is the $i$-th cluster. $c_{i}$ is the centroid of cluster $C_{i}$. $x$ is the feature of the target node.
    (7) $Cosine(x,y)=\frac{\sum_{i=1}^{n}x_{i}y_{i}}{\sqrt{\sum_{i=1}^{n}x_{i}^{2}}\sqrt{\sum_{i=1}^{n}y_{i}^{2}}}.$
B.2. Implementation details
We implement our ArchRAG in Python, while C-HNSW is implemented in C++ and provides a Python interface for integration. We implement C-HNSW using the FAISS framework and employ the inner product metric to measure the proximity between two vectors. All methods utilize 10 concurrent LLM calls, and to maintain consistency, other parallel computations in the method, such as embedding calculations, also use 10 concurrent threads. All the experiments were conducted on a Linux operating system running on a machine with an Intel Xeon 2.0GHz CPU, 1024GB of memory, and 8 NVIDIA GeForce RTX A5000 GPUs, each with 24 GB of memory. Figure 10 demonstrates the prompt used in Adaptive filtering-based generation. Please refer to our repository (anonymous.4open.science/r/H-CAR-AG-1E0B) to view the detailed prompts.
Prompt for Adaptive filtering-based generation Filger Prompt: # Role You are a helpful assistant responding to questions about data in the tables provided. # Goal Generate a response consisting of a list of key points that respond to the user’s question, summarizing all relevant information in the input data tables. You should use the data provided in the data tables below as the primary context for generating the response. If you don’t know the answer or if the input data tables do not contain sufficient information to provide an answer, just say so. Do not make anything up. Each key point in the response should have the following element: Description: A comprehensive description of the point. Importance Score: An integer score between 0-100 that indicates how important the point is in answering the user’s question. An ’I don’t know’ type of response should have a score of 0. The response should be JSON formatted as follows:
    "points": [
        {"description": "Description of point 1 [Data: Reports (report ids)]", "score": score_value},
        {"description": "Description of point 2 [Data: Reports (report ids)]", "score": score_value},
# User Question
    {user_query}
# Data tables
    {context_data}
Output:
Merge Prompt: # Role You are a helpful assistant responding to questions and may use the provided data as a reference. # Goal You should incorporate insights from all the reports from multiple analysts who focused on different parts of the dataset to support your answer. Please note that the provided information may contain inaccuracies or be unrelated. If the provided information does not address the question, please respond using what you know. namely, A response that utilizes the provided information, ensuring that all irrelevant details from the analysts’ reports are removed. A response to the user’s query based on your existing knowledge when ¡Analyst Reports¿ is empty. The final response should merge the relevant information into a comprehensive answer that clearly explains all key points and implications, tailored to the appropriate response length and format. Note that the analysts’ reports provided below are ranked in the descending order of importance. Do not include information where the supporting evidence for it is not provided. # Target response length and format
    {response_format}
# User Question
    {user_query}
# Analyst Reports
    {report_data}
Output:
Figure 10. The prompt for adaptive filtering-based generation.